OpenStudy (anonymous):

Can some one explain to me what kinda problem this iz a citys population can be modeld by this equation p=130(1.03)^x where x is the number of years after january 1, 2000 for what value f x dose this model predict that the population of the city will be 170,000

OpenStudy (anonymous):

This is what we call an exponential graph. If you have a graphing calculator, you can plug in this equation. Then look for the place where p = 170,000 If you don't have a graphing calculator, this can still be solved pretty easily. Set the equation to: 170000=130(1.03)^x This becomes 170000=133.9^x Take the log of both sides. log 170000=log 133.9^x If you do this, your answer is: x = 2.45932701188 So after 2.46 years, the population will be at 170,000.

OpenStudy (anonymous):

Thanxx